### v1

# Magnetic effects in the Hadron Resonance Gas

Abstract
We discuss the modeling of the hadronic phase of QCD at finite magnetic field in the framework of hadron resonance gas (HRG). We focus on the statistical description of particle yields that include contribution from resonance decays. We demonstrate that the swift increase in the number of protons with magnetic field predicted in the HRG is due to the ill-defined description of higher-spin states. We discuss fluctuations of conserved charges and show that at present the qualitative comparison of the model predictions with the Lattice QCD data should be treated with care. We also discuss the principle of detailed balance which allows to study the magnetic field dependence of neutral resonances.
I Introduction
Magnetic fields generated in non-central heavy-ion collisions (HICs) are among the strongest in the universe [1, 2, 3]. Since quarks and antiquarks are electrically charged and influenced by magnetic field, understanding the interplay between the magnetic field and the strong interactions governed by quantum chromodynamics (QCD) is crucial for properly comprehending non-central heavy-ion collisions. Understanding the impact of an extremely strong magnetic field on the strongly interacting matter is also relevant for studying magnetars [4] and the early universe [5]. The magnetic field induces a variety of interesting phenomena in the QCD matter, such as anomaly-induced transport phenomena e.g., Chiral Magnetic Effect (CME) [6] and Chiral Vortical Effect (CVE) [7], charged $\rho$ meson condensation [8], modification of QCD Debye mass [9], QCD equation of state (EoS) [10, 9, 11, 12], dilepton production rate [13, 14].
Investigation of QCD in the background of magnetic fields is also interesting from the theoretical standpoint since it can be simulated from the first principle lattice QCD (LQCD) methods even for large fields in contrast to finite baryon chemical potential where the simulations are marred by the numerical sign problem. LQCD simulations indicate a nontrivial interplay between the magnetic field and strong interactions, such as magnetic catalysis at low temperatures [15, 16], and inverse magnetic catalysis [17, 18, 19] which influences the QCD phase diagram. These effects have also been studied widely using several other formalisms viz., Nambu-Jona-Lasinio (NJL) model, Polyakov improved NJL (PNJL) model [20], quark-meson (QM) model [20], gauge-gravity duality [21]. However, the latter effect is not captured in commonly used chiral models [22]. It can, nevertheless, be generated by considering additional in-medium effects (see e.g. [23]). It has also been found that the pseudocritical temperature of the QCD crossover decreases with the magnetic field (see, e.g., Refs. [24, 25, 18]), with the possible existence of the critical point at large field intensity [25, 26]. For recent reviews of QCD in external magnetic field, see [22, 27, 28, 29]. Recently, there has been significant attention to the fluctuations of conserved charges in the magnetic field background using LQCD methods [30, 31, 12, 32]. It has been observed that a finite magnetic field has a nontrivial effect on these observables. They are of direct phenomenological interest because they are sensitive to in-medium degrees of freedom and can be measured experimentally.
The hadron resonance gas (HRG) model has been successful in describing the LQCD data on the EoS and fluctuations of conserved charges at vanishing magnetic field [33, 34] as well as the hadron yields from heavy-ion collisions [35]. In [36], the HRG model has been generalized to include the effects of finite magnetic field. The model has been used to study the impact of magnetic field on the QCD EoS [36, 10], fluctuations of conserved charges [37, 38, 31, 39] as well as other effects, such as transport coefficients [40, 41]. The behavior of spin-00 and spin-$1/2$ particles in an external magnetic field is well understood and involves the Landau quantization in the directions perpendicular to the magnetic field [42]. However, when the same prescription is applied to higher-spin particles, it leads to imaginary dispersion relations at strong fields as well as instabilities in the pressure [36]. This suggests that the treatment of higher-spin particles in the HRG model is doubtful and requires further investigation.
In this work, we study the fluctuations of conserved charges and particle yields at a finite magnetic field using the HRG model. We find that the magnetic field enhances the baryon yields, while $\pi^{\pm}$ and $K^{\pm}$ yields are weakly affected. With our choice of resonances included in HRG, we find that the model underestimates the LQCD baryon number fluctuations at a weak magnetic field while overestimating them at large $eB$, the latter originating from the divergences induced by higher-spin resonances. We also gauge the impact of the magnetic field on neutral resonances by implementing certain detailed balance conditions. We compare the proton and pion yields obtained using HRG and our detailed balance requirement and observe a substantial difference, indicating the need for further research on the impact of strong magnetic field on resonance properties.
This paper is organized as follows: In Sec. II we introduce the HRG thermodynamics at finite temperature and magnetic field and show the numerical results. We discuss the effect of magnetic field on particle densities and yields, and fluctuations of conserved charges in conjunction with the LQCD data. In Sec. III, we show how the external magnetic field affects neutral particle densities considering detailed balance conditions. Finally, Sec. IV is devoted to summary and conclusions.
II Hadronic Matter in finite magnetic field
In a phenomenological description of hadronic matter, one needs to identify the relevant degrees of freedom and their interactions. In the confined phase of QCD, the medium is composed of hadrons and their resonances. In its simplest version, the hadron resonance gas (HRG) model assumes that the constituents of the medium are independent and point-like [43]. This effectively neglects their widths and interactions. Consequently, the pressure in the HRG model is approximated by the sum over partial pressures of hadrons and their resonances, treated as noninteracting particles,
$P=\sum\limits_{i}P_{i}\rm,$ (1)
where $i$ goes through all strange and non-strange hadrons and their resonances listed in the Particle Data Group summary tables [44]#1#1#1In this work, we include established mesons and baryons with three- and four-star rating.. We note that the thermodynamic pressure $P$ contains all the relevant information about the medium through the mass and quantum numbers of hadrons. Thus, it allows for the study of different thermodynamic observables, including particle numbers and fluctuations of conserved charges.
The partial pressures in Eq. (1) are given as
$P_{i}=\pm\gamma_{i}T\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\;\log\left(1\pm f_{i}\right)\textrm{,}$ (2)
where $\gamma_{i}$ is the spin degeneracy factor and
$f_{i}=\frac{1}{e^{\beta(\epsilon_{i}-\mu_{i})}\mp 1}$ (3)
is the distribution function, $\epsilon_{i}=\sqrt{\vec{p}^{2}+m_{i}^{2}}$ is the dispersion relation and $\mu_{i}=B_{i}\mu_{B}+Q_{i}\mu_{Q}+S_{i}\mu_{S}$ is the chemical potential of the particle. The upper (lower) signs in Eqs. (2) and (3) refer to fermions (bosons).
Figure 1: Particle densities normalized to densities at vanishing magnetic field at $T=0.155~{}$GeV as functions of magnetic field $eB$.
In the presence of a constant external magnetic field $eB$ pointing along the $z$ direction, the system undergoes Landau quantization in the $xy$ plane [22, 27]. Consequently, the dispersion relation for a charged particle $(Q>0)$ becomes
$\epsilon=\sqrt{p_{z}^{2}+m^{2}+2|Q|B\left(l+\frac{1}{2}-s_{z}\right)}\textrm{,}$ (4)
where $s_{z}$ is the $z$-component of the particle’s spin and $l\in\{0,1,2,\ldots\}$ numbers the Landau levels. We note that the dispersion relation in Eq. (4) is exact for structureless spin-00 and spin-$1/2$ particles [22]. It has also been found that spin-$3/2$ particles, governed by the Rarita-Schwinger equation, exhibit a non-causal behavior in the presence of external magnetic field [45, 46, 47], which can be fixed by adding non-minimal interaction term to the Lagrangian [48, 49] and end up with a similar form of dispersion relation depicted at Eq. (4).
We stress that Eq. (4) neglects the compositeness of the state. Theoretical issues can arise when the scale of magnetic field resolves the structure of hadron states, i.e., $\sqrt{|Q|B}>m_{\pi}$. The problem becomes more apparent for higher spin states, where the mass of particles can be substantially reduced and eventually leads to a complex dispersion relation (signaling instability). Resolving these issues would require a more sophisticated treatment of the decay dynamics [50, 51].
The phase-space integral of a charged particle acquires the following form [52, 53]
$\gamma\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\rightarrow\frac{|Q|B}{2\pi^{2}}\sum_{s_{z}}\sum_{l=0}^{\infty}\int\limits_{0}^{\infty}\mathrm{d}p_{z}\textrm{,}$ (5)
where $\gamma$ is the spin degeneracy factor and $Q$ is the electric charge of the particle. As a result, the thermal part of the pressure becomes
$P_{i}=\pm T\frac{|Q_{i}|B}{2\pi^{2}}\sum_{s_{z}}\sum_{l=0}^{\infty}\int\limits_{0}^{\infty}\mathrm{d}p_{z}\;\log\left(1\pm f_{i}\right)\,.$ (6)
Figure 2: Particle yields normalized to the yields at vanishing magnetic field at $T=0.155~{}$GeV as functions of the magnetic field $eB$.
We note that the pressure of a finite magnetic field contributes to the bulk thermodynamic quantities, and the vacuum part of the thermodynamic pressure must be regularized [36]. In this work, we focus on the densities and fluctuations of conserved charges which do not depend on the vacuum part of the pressure. Thus, we omit it in our considerations. We also note that the magnetic field does not distinguish between positive and negative electric charges (cf. Eqs. (4) and (6) ).
The particle density is defined as
$n_{i}=\frac{\langle N_{i}\rangle}{V}=\frac{\partial P_{i}}{\partial\mu_{i}}\Bigg{|}_{T}\rm,$ (7)
where $\langle N_{i}\rangle$ is an average number of particle $i$ in volume $V$. To quantify the change of the medium composition due to the presence of a finite magnetic field, we study the normalized densities of selected hadrons at fixed temperature $T=0.155~{}\rm GeV$. This is depicted in Fig. 1. The densities of $p$, $\Sigma^{+}$, and $\Xi^{-}$ generally increase with the magnetic field. This is attributed to the lowest Landau level being magnetic field-independent and hence its contribution grows linearly with $eB$ (cf. Eqs. (4) and (6)). In contrast to this, the densities of $\pi^{-}$ and $K^{-}$ are exponentially suppressed because in this case, all Landau levels increase with magnetic field. We note that there is a clear mass ordering for spin-$1/2$ baryons and spin-00 mesons. The exceptions to mass ordering are the higher-spin particles, $\rho^{+}$ and $\Omega^{-}$, whose increase is much greater than for other particle species. This swift increase is caused by the contribution of the lowest Landau level being negative for these particles, see Eq.(4). As mentioned, the extrapolation of Eq. (4) to higher-spin states requires the assumption of structureless hadrons, which can be questionable depending on the size of magnetic field and resonance structure of the state in question.
Figure 3: Particle yields normalized to the $\pi^{+}$ yield at $T=0.155~{}$GeV as functions of magnetic field $eB$.
The HRG model can be used to determine the hadron yields, assuming thermal and chemical equilibrium between all stable hadrons and resonances (see, e.g., [35, 54, 55]). This is done by accounting for the resonance decays into lighter particles. The average number $\langle N_{i}\rangle$ of particle $i$ in volume $V$ is given as [43]
$\langle N_{i}\rangle=\langle N^{\rm th}_{i}\rangle+\sum_{r\rightarrow i}\Gamma_{r\rightarrow i}\langle N_{r}\rangle\rm,$ (8)
where $N^{\rm th}_{i}=Vn_{i}$ is the average thermal number defined through Eq. (7), the sum in the second term is overall decays with particle $i$ in the final state, and $\Gamma_{r\rightarrow i}$ is the branching ratio of the resonance $r$ to particle $i$.
In Fig. 2, we show the particle yields normalized to the yields at the vanishing magnetic field. At small values of $eB$, the $\pi^{+}$ and $K^{+}$ yields decrease with increasing $eB$, but at a stronger magnetic field, their yields increase with increasing $eB$. Nevertheless, at $eB\lesssim 0.2~{}\rm GeV^{2}$, $\pi^{+}$ and $K^{+}$ yields change only a few percent. However, at the same values of $eB$, baryon yields increase substantially. The largest increase is seen for protons, due to decaying $\Delta$ resonances. We warn the reader again about the uncertainties related in the description of high spin ($s\geq 1$) states and note that the weak dependency of $\Sigma$ and $\Xi$ yields on the magnetic field may be a result of the way fewer known strange than non-strange resonances.
In the literature, the observed particle yields are often discussed in terms of particle ratios instead of actual yields to avoid dependency on system size. To facilitate the discussion in terms of particle ratios, we show the particle yields normalized to the yield of $\pi^{+}$ in Fig. 3. Again, the prominent feature is the substantial increase of the $p/\pi$ ratio with the increasing strength of the magnetic field. We note that even if the $\Omega^{+}$ yield increases way faster than the pion yield in Fig. 2, it is still so tiny compared to the pion yield that the change in the ratio cannot be seen at the scale of Fig. 3.
Figure 4: Net-baryon number susceptibility for $T=0.145$ (red) and 0.1550.1550.155~{}0.155GeV (blue) as a function of magnetic field $eB$.
Now we turn to the fluctuations of conserved charges. In a thermal medium, the second-order fluctuations and correlations of conserved charges are quantified by the generalized susceptibilities,
$\chi_{ij}=\frac{\partial^{2}(P/T^{4})}{\partial(\mu_{i}/T)\partial(\mu_{j}/T)}\Bigg{|}_{T}.$ (9)
where $i,j\in\{B,~{}Q,~{}S\}$. The susceptibilities are observables sensitive to the quantum numbers of medium constituents. Thus, they can be used to identify the contributions of different species of particles to QCD thermodynamics [56, 57].
For a sufficiently small net-baryon density, the EoS can be expanded in the Taylor series. Under the assumption of strangeness neutrality, the pressure is expanded as [32]
$\frac{P}{T^{4}}\approx c_{0}+c_{2}\left(\frac{\mu_{B}}{T}\right)^{2}+\mathcal{O}\left(\left(\frac{\mu_{B}}{T}\right)^{4}\right)\rm,$ (10)
where $c_{0}$ is the pressure at vanishing chemical potentials and $c_{2}$ is a mixture of second-order susceptibilities
$c_{2}=\frac{1}{2}\chi_{BB}+\frac{1}{3}\chi_{BS}+\frac{1}{18}\chi_{SS}\rm.$ (11)
We note that the coefficient $c_{2}$ is evaluated at vanishing chemical potentials. At a sufficiently small net-baryon density, the strangeness neutrality condition can be approximated by $\mu_{S}=0$. Thus, one can expect that $\chi_{BB}\approx 2c_{2}$, as is found to be the case in the HRG model. This allows us to compile the LQCD data on $\chi_{BB}$ [31] and $c_{2}$ [32] in a single figure. They are shown in Fig. 4 for $T=0.145$ and 0.1550.1550.155~{}0.155GeV. We find that the HRG model systematically underestimates the LQCD data for $eB\leq 0.16~{}\rm GeV^{2}$ at both temperatures. We note that this is in contrast to the HRG model with an augmented list of resonances which includes states predicted by relativistic quark models. Such a HRG model reproduces $\chi_{BB}$ at vanishing magnetic field [34], but, as the magnetic field increases, both the HRG model with established states only and the model with all the predicted states, start to overestimate the LQCD results. The calculated susceptibility increases rapidly and drastically overshoots the LQCD data at $eB=0.3~{}\rm GeV^{2}$. The contribution of baryons with $s\geq 3/2$ ranges from $75\%$ at vanishing $eB$ to $95\%$ at $eB=0.3~{}\rm GeV^{2}$. This result exposes the weakness of the current treatment of the higher-spin resonances.
Ultimately, to fully identify the contributions of higher-spin states to thermodynamic observables their internal structure should be taken into account. This, however, requires a dynamical treatment of resonances, e.g. the S-matrix formulation of statistical mechanics [58, 59, 60], suitably extended to the presence of $eB$, and is currently under development. As a temporary attempt, we consider the detailed balance approach to neutral particles (see the next Section).
Our major observation is that the current uncertain treatment of high-spin particles in a finite magnetic field prevents direct comparison of the HRG model with the experimental data from HIC and results obtained in LQCD simulations.
Figure 5: $\rho^{0}$ and $\Delta^{0}$ densities when the detailed balance is required, normalized to densities at $eB=0$, at $T=0.1$ and 0.155 GeV as functions of the magnetic field.
III Detailed Balance
As mentioned, we assume neutral particles to have zero magnetic moments. Consequently, they are not affected by the magnetic field at all. However, in chemically equilibrated HRG detailed balance prevails, e.g., the decay rate of rho mesons is equal to the scattering rate of pions forming $\rho$ resonances. The scattering rate depends on the densities and, as seen in Fig. 1, pion density decreases with the magnetic field. Thus, the $\rho^{0}$ density may depend on the magnetic field after all.
It is known that a weak magnetic field causes a negligibly small change in the width of $\rho^{0}$ [50]. Thus in the presence of a magnetic field, the $\rho^{0}$ decay rate changes only if the $\rho^{0}$ density changes, and the change in rate is proportional to the change in density.
Unfortunately, we cannot evaluate the scattering rate of pions using the conventional kinetic theory calculation, since particles in magnetic field do not have well-defined momenta. To gain insight into how the detailed balance requirement could affect the densities of neutral resonances, we make the bold assumption that the magnetic field affects the pion scattering rate (and thus $\rho^{0}$ production rate) only by changing their densities. In other words, at fixed temperature,
$\Gamma(\pi^{+}\pi^{-}\rightarrow\rho^{0})=\frac{n_{\pi^{+}}(B)\,n_{\pi^{-}}(B)}{n^{eq}_{\pi^{+}}\,n^{eq}_{\pi^{-}}}\Gamma^{eq}(\pi^{+}\pi^{-}\rightarrow\rho^{0}),$ (12)
where $\Gamma(\pi^{+}\pi^{-}\rightarrow\rho^{0})$ is the production rate of $\rho^{0}$ in the presence of a magnetic field, $n_{\pi^{\pm}}(B)$ are the pion densities in magnetic field, $n^{eq}_{\pi^{\pm}}$ are the corresponding pion densities when $eB=0$, and $\Gamma^{eq}(\pi^{+}\pi^{-}\rightarrow\rho^{0})$ the $\rho^{0}$ production rate in equilibrated HRG when there is no magnetic field.
Figure 6: Proton and pion yields (top panel) and their ratio (bottom panel) as functions of magnetic field $eB$ calculated using the conventional HRG in magnetic field approach (HRG) and using the requirement of detailed balance discussed in Sect. III (this work). The results are for $T=0.155~{}\rm GeV$ and are normalized to their values at vanishing magnetic field.
We can evaluate an effective pion chemical potential such that
$n_{\pi}(T,\mu_{\pi},B=0)=n_{\pi}(T,B).$ (13)
With the above-mentioned considerations, detailed balance requires that $\rho^{0}$ obtains a chemical potential, which is a sum of pion chemical potentials: $\mu_{\rho^{0}}=\mu_{\pi^{+}}+\mu_{\pi^{-}}=2\mu_{\pi^{+}}$, since both $\pi^{+}$ and $\pi^{-}$ are affected the same way by the magnetic field, and thus their effective chemical potentials are equal. We can repeat similar arguments for $\Delta^{0}$ resonance, and obtain a requirement that it also develops an effective chemical potential in magnetic field: $\mu_{\Delta^{0}}=(\mu_{p}+\mu_{\pi^{+}})/3$.
We have evaluated the $\rho^{0}$ and $\Delta^{0}$ densities corresponding to these chemical potentials, and show them scaled with their equilibrium densities in Fig. 5. Unlike the expectation that neutral particles should not be affected by the magnetic field, both $\rho^{0}$ and $\Delta^{0}$ densities decrease with increasing strength of magnetic field, the effect being the stronger the colder the system. We note here that our assumption about the production rate is extremely naive, but the $\rho^{0}$ and $\Delta^{0}$ densities staying independent of the magnetic field would require very careful fine-tuning of the pion-pion and pion-nucleon interactions.
Detailed balance must prevail not only between neutral resonances and their daughters but also between charged resonances and their daughters. To gauge the overall uncertainty in the chemical composition of HRG in magnetic field, we assume that $\Delta$ and $\rho$ densities are not given by Eqs. 47, but all charged states of $\Delta$ and $\rho$ gain effective chemical potentials similar to $\Delta^{0}$ and $\rho^{0}$. We keep all the other densities unchanged and evaluate the pion and proton yields after decays. The resulting yields and their ratios are shown in Fig. 6. As expected, the reduction in $\rho$ and $\Delta$ densities leads to reduced yields, and since the reduction in $\rho$ is larger than in $\Delta$, the proton to pion ratio is larger.
In the literature, the centrality dependence of proton to pion ratio has been suggested as a measure of the magnetic field in heavy-ion collisions, with the assumption that the field is negligible in most central collisions, and largest in most peripheral. The ALICE collaboration has measured the $p/\pi$ ratio in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\rm TeV$ [61], and reported that the ratio increases by a factor $\sim 1.3$ from central to peripheral collisions. In the lower panel of Fig. 6, we have included dots on the $p/\pi$ curves indicating where the ratio has increased by 30% compared to its value at vanishing magnetic field. We see that our assumptions about the chemical composition of HRG lead to similar $\sim 30$% uncertainty in the allowed strength of the magnetic field.
IV Conclusions
In this work, we examined the effect of finite magnetic field on the thermal properties of hadronic medium. We modeled the hadronic phase of QCD with the hadron resonance gas (HRG) model. For spin-00 particles the densities decrease with magnetic field, while the densities of spin-$1/2$ show the opposite behavior which is due to a different structure of the lowest Landau level. Densities of higher-spin resonances strongly increase with magnetic field which is caused by a singular behavior of the dispersion relation (cf. Eq. (4)) applied to higher-spin particles. Since these densities contribute to final production yields of proton and pions after including the resonance decays, the latter also diverge at sufficiently large fields. This may be problematic for a reliable estimation of magnetic field from non-central HICs based on observed particle yields. As an example, we compared the magnetic field predicted using the HRG and the detailed balance conditions and found a substantial difference between both predictions. Our results suggest that the interpretation of the magnetic field dependence of particle yields in the HRG model, although feasible, should be treated with great care.
We have also studied the second-order fluctuations of conserved charges in conjunction with the recent LQCD data. We find that for our choice of the resonance content of HRG, the LQCD data is underestimated by model results at small fields, while the singular behavior at larger $eB$ leads to overestimation of the data. Although considering hypothetical states improves the description of LQCD data at low $eB$ [31], it makes the breakdown in strong fields even worse.
Therefore, for a reliable comparison between LQCD results and experimental data, it would be crucial to treat the dynamical structure of resonances, especially those with higher spins and neutral states. As the first step, we explored the detailed balance approach and obtained promising results. Another possibility is to generalize the S-matrix approach to finite magnetic field which would allow us to take into account the dynamics of hadrons via the scattering data. Ultimately, non-perturbative approaches capable of describing bound states should be employed to study the dynamics of hadrons in an external magnetic field. Many of these interesting points will be pursued in future research.
Acknowledgments
This work is supported by the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science (P.H, M.M. and B.K.). M.S. acknowledges the financial support of the Polish National Science Center (NCN) under the Preludium grant 2020/37/N/ST2/00367. This work is supported partly by the Polish National Science Centre (NCN) under OPUS Grant No. 2022/45/B/ST2/01527 (C.S., K.R. and P.M.L.). K.R. also acknowledges the support of the Polish Ministry of Science and Higher Education. The work of C.S. was supported in part by the World Premier International Research Center Initiative (WPI) under MEXT, Japan.